# Role of activity and dissipation in achieving precise beating in cilia: Insights from the rower model

### v1

@@ III.4 Energy dissipation rate
Achieving high precision and the associated energy budget has been a key recent focus in biophysics research [84, 85, 86, 87]. The thermodynamic uncertainty relation suggests the following inequality for the precision quality factor: ${\cal Q}\leq{\cal I}$ where ${\cal I}=\frac{q_{avg}}{2D\gamma}$, where $q_{avg}$ is the dissipated heat over the driving time [70], requiring a higher energy cost for a higher precision. However, as shown in  [87], a cyclic external protocol driving a system to a periodic steady state can achieve high precision with an arbitrarily small energy budget.
Given this context, we ask about the energy budget required to achieve precision ${\cal Q}$ in the rower model. The energy input per cycle, through two switching events, is $2k\mu\mathcal{A}$, which is fully dissipated. Due to symmetry, the dissipation rates for $\sigma=1$ and $\sigma=-1$ strokes are identical. The average energy dissipation rate during the movement from $-\mathcal{A}$ to $\mathcal{A}$ over the first passage time $\mathcal{T}$ is given by:
$\dot{q}_{avg}$ $=$ $\left\langle\frac{1}{{\cal T}}\int_{0}^{\cal T}{-}\frac{dV(x,\sigma=1)}{dx}\dot{x}dt\right\rangle$ (15)
$\approx$ $\frac{k\mu\mathcal{A}}{\langle\mathcal{T}\rangle}=\frac{k^{2}\mu^{2}\mathcal{A}_{s}}{2\gamma\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right)}.$
@@
Despite energy dissipation decreasing monotonically with scaled amplitude ${\cal A}_{s}$, oscillation precision ${\cal Q}$ reaches its maximum at ${\cal A}_{s}=0.518$, independent of other parameters.
Figure 5: Energy dissipation rate $\dot{q}_{avg}$ versus scaled amplitude $\mathcal{A}_{s}$. Symbols denote simulation results, and the line represents the analytical result (Eq.(15)). Parameters: $\mu=2\ \mu$m, $k=4\ pN/\mu$m, and $D=0.02$.
As pointed out, the energy dissipated over an average period $\langle\cal T\rangle$, $q_{avg}=\frac{1}{2}k\mu^{2}{\cal A}_{s}$ and thus ${\cal I}=\frac{k\mu^{2}}{4D\gamma}{\cal A}_{s}$, although ${\cal Q}=\frac{k\mu^{2}}{D\gamma}g({\cal A}_{s})$. This leads to the relation
@@
As the expression shows, $0\leq\frac{{\cal Q}}{{\cal I}}\leq 1$, with the bounds defined by ${\cal A}_{s}=1$ and ${\cal A}_{s}=0$. This is consistent with the thermodynamic uncertainty relation, ${\cal Q}/{\cal I}\leq 1$. For any useful oscillation, ${\cal A}_{s}>0$ is required, enforcing a stricter constraint, ${\cal Q}/{\cal I}<1$. The exact expression for ${\cal Q}/{\cal I}$ in Eq.(17) is obtained within SNA but the upper bound to the ratio is more general.
IV Comparison with the rotor model
@@
where $\xi(t)$ is a Gaussian white noise with $\langle\xi\rangle=0$ and $\langle\xi(t)\xi(0)\rangle=2D\delta(t)$. In rotor synchronization studies, the force $F$ must vary with $\theta$ [49, 51, 50, 53, 52, 47, 61], but for simplicity, here we assume $F=F_{0}$, a constant. The rotor model pumps energy continuously along the trajectory, unlike the rower model, which pumps at switching points. The rotor has a single length scale $R$, while the rower model involves two: $\mathcal{A}$ and $\mu$. The energy per cycle is $2\pi F_{0}R$ for the rotor, and $2k\mu\mathcal{A}$ for the rower.
Using SNA, the mean and variance for the time $\mathcal{T}$ to reach $\theta=2\pi$ from $\theta=0$ are given by
@@ V Discussions and conclusion
The precision of oscillations is typically quantified by the quality factor $Q$, the ratio of coherence time to period. As we have shown, this is equivalent to another precision quality factor ${\cal Q}$, defined as the ratio of mean current to current fluctuation. This quantity is used in stochastic thermodynamics to establish a thermodynamic uncertainty relation, setting an upper limit based on dissipation. We use the SNA to express this quantity in terms of the relative variance in first passage time (FPT), $CV^{2}(\mathcal{T})$. In the rower model, the first passage time (FPT) $\mathcal{T}$ is the time for the beating to move from $-\mathcal{A}$ to $\mathcal{A}$ in a harmonic potential with stiffness $k$ centered at $\mu/2$ within a viscous medium with drag coefficient $\gamma$. The SNA provides a simple formula for $CV^{2}(\mathcal{T})$, matching numerical results closely. In the absence of a closed-form expression, the exact integral form of $CV^{2}(\mathcal{T})$ provides limited analytical clarity [81, 77], whereas the approximate SNA offers more profound insights. We found that $CV^{2}(\mathcal{T})$ is a function of scaled amplitude $\mathcal{A}_{s}=2\mathcal{A}/\mu$, with a scaling factor $k\mu^{2}/D\gamma$, representing the ratio of active energy scale to effective temperature. The SNA allows us to derive a closed-form expression of the scaling function $g({\cal A}_{s})$ which reaches a maximum at ${\cal A}_{s}=0.518$, irrespective of other parameters. Our analytical expression for the ratio of precision quality factor ${\cal Q}$ and energy dissipation per cycle $q_{avg}$ gives a thermodynamic uncertainty relation ${\cal Q}\,(2D\gamma/q_{avg})\leq 1$.
@@
DC and SD designed and supervised the research. SG performed the numerical calculations. All authors discussed the results and contributed to the manuscript.
@@
The data are available from the authors upon reasonable request.
@@ Acknowledgements
VI Appendix
VI.1 Small noise approximation for FPT statistics
The displacement of the bead $x$ is a random variable that fluctuates in time $t$ and its dynamics is given by the Langevin equation presented in the main text (Eq. 2). For a harmonic potential, it is possible to solve dynamics for the mean and variance of displacement. In the first passage problem for $x(t)$ to reach a final position $x_{f}$ for the first time starting from the initial position $x_{0}$, the time $t$ becomes a random variable. For many random processes including random motions in harmonic potentials, exact solutions of FPT statistics exist. Mathematical calculations for FPT are generally involved, and the analytical expressions are often complicated, which hinders simple analytical insights.
If the fluctuations in $x$ are small and the dynamics of the mean and variance are known, then one can obtain an approximated formula for noise in FPT using the latter. This SNA has been used in various contexts and is quite useful for theoretical understanding, especially for complex dynamics where exact solutions are impossible.
Let us assume that $t$ fluctuates around the mean $\langle t\rangle$, i.e., $t=\langle t\rangle+\delta t$, where the fluctuation $\delta t$ is small. Then, ignoring higher-order terms in the Taylor expansion of $x(t)$, we get
$x(t)$$\approx$ $\langle x(t)\rangle+\delta t\frac{dx}{dt}\Big{|}_{t=\langle t\rangle}.$ (21)
Taking the square of both sides of Eq. 21 and then taking the expected value gives
$\langle\delta x^{2}\rangle$$\approx$ $\langle\delta t^{2}\rangle\left(\frac{d\langle x\rangle}{dt}\right)^{2}\Big{|}_{t=\langle t\rangle},$ (22)
where, $\delta x=x(t)-\langle x(t)\rangle$. In terms of the squared coefficient of variation or relative variance, the Eq. 22 can be written as
$CV^{2}(t)=\frac{\langle\delta t^{2}\rangle}{\langle t\rangle^{2}}$$\approx$ $\frac{\langle\delta x^{2}\rangle}{\langle t\rangle^{2}}\left(\frac{d\langle x\rangle}{dt}\right)^{-2}\Big{|}_{t=\langle t\rangle},$ (23)
$\approx$ $CV^{2}(x)\left(\frac{\langle t\rangle}{\langle x\rangle}\frac{d\langle x\rangle}{dt}\right)^{-2}\Bigg{|}_{t=\langle t\rangle}.$
We are interested in the dynamics of mean and variance in $x$ for $\sigma=1$ potential. Solving Eq. 2 with the initial condition $x_{0}=-\mathcal{A}$ we obtain,
$\langle x(t)\rangle{=}-\mathcal{A}\exp\left(-\frac{kt}{\gamma}\right){+}\frac{\mu}{2}\left[1{-}\exp\left(-\frac{kt}{\gamma}\right)\right],$ (24)
$\langle x(t)^{2}\rangle-\langle x(t)\rangle^{2}{=}\frac{D\gamma}{k}\left[1{-}\exp\left(-\frac{2kt}{\gamma}\right)\right].$ (25)
The mean FPT time to reach $\mathcal{A}$ from $-\mathcal{A}$ can be obtained by solving Eq. 24 for time and given by
$t=\langle\mathcal{T}\rangle$ $=$ $\frac{\gamma}{k}\log\left(\frac{\mu+2\mathcal{A}}{\mu-2\mathcal{A}}\right).$ (26)
$=$ $\frac{\gamma}{k}\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right),$
where $\mathcal{A}_{s}=2\mathcal{A}/\mu$. The mean FPT is nothing but half of the period. Using Eqs. 24, 25, and 26, we finally derive the expression for noise in FPT as
$CV^{2}(\mathcal{T})$ $=$ $\frac{D\gamma}{k\mu^{2}}\frac{16\mathcal{A}_{s}}{\left((1-\mathcal{A}_{s}^{2})\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right)\right)^{2}}.$ (27)

### v2

@@ III.4 Energy dissipation rate
Achieving high precision and the associated energy budget has been a key recent focus in biophysics research [16, 68, 47, 7]. The thermodynamic uncertainty relation states the following inequality for the precision quality factor: $\cal Q$$\leq{\cal I}$ where ${\cal I}=\frac{q_{avg}}{2D\gamma}$, where $q_{avg}$ is the dissipated heat over the driving time [6], requiring a higher energy cost for a higher precision. However, as shown in  [7], a cyclic external protocol driving a system to a periodic steady state can achieve high precision with an arbitrarily small energy budget.
Given this context, we ask about the energy budget required to achieve precision $\cal Q$} in the rower model. The energy input per cycle, through two switching events, is $2k\mu\mathcal{A}$, which is fully dissipated. Due to symmetry, the dissipation rates for $\sigma=1$ and $\sigma=-1$ strokes are identical. The average energy dissipation rate during the movement from $\mathcal{A}$ to $\mathcal{A}$ over the first passage time $\tau_{d}\mathcal{T}$ is given by:
$\dot{q}_{avg}$ $=$ $\left\langle\frac{1}{\tau_{d}{\cal T}}\int_{0}^{\tau_{d}\cal T}{-}\frac{dV(x,\sigma=1)}{dx}\dot{x}dt\right\rangle$ (18)
$\approx$ $\frac{k\mu\mathcal{A}}{\tau_{d}\langle\mathcal{T}\rangle}=\frac{k^{2}\mu^{2}\mathcal{A}_{s}}{2\gamma\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right)}.$
@@
Despite energy dissipation rate decreasing monotonically with scaled amplitude ${\cal A}_{s}$, oscillation precision $\cal Q$} reaches its maximum at ${\cal A}_{s}=0.518$, independent of other parameters.
[…]
As pointed out, the energy dissipated over an average period $\tau_{d}\langle\cal T\rangle$, $q_{avg}=\frac{1}{2}k\mu^{2}{\cal A}_{s}$ and thus ${\cal I}=\frac{k\mu^{2}}{4D\gamma}{\cal A}_{s}$, although ${\cal Q}=\frac{k\mu^{2}}{D\gamma}g({\cal A}_{s})$. This leads to the relation
@@
[…]
IV Precision in asymmetric beating of cilia
Thus far, our analysis has assumed symmetric activity and dissipation in both beating directions. In reality, however, the forward and reverse beats of cilia are generally asymmetric. In this section, we examine how such asymmetry affects oscillation precision by studying an asymmetric rower model. The asymmetry may arise from differences in drag forces or from an asymmetric potential well [46, 87].
Here, we analyze the first-passage-time (FPT) statistics for the two branches separately. Let $\mathcal{T}_{\pm 1}$ denote the dimensionless first-passage time for $\sigma=\pm 1$. In the steady state, the total period is $\mathcal{T}=\mathcal{T}_{+1}+\mathcal{T}_{-1}$, and the fluctuations in the period can be written as,
$CV^{2}(\mathcal{T}){=}\frac{CV^{2}(\mathcal{T}_{+1})\langle\mathcal{T}_{+1}\rangle^{2}{+}\,CV^{2}(\mathcal{T}_{-1})\langle\mathcal{T}_{-1}\rangle^{2}}{\langle\mathcal{T}_{+1}{+}\mathcal{T}_{-1}\rangle^{2}},$ (21)
where $CV^{2}(\mathcal{T}_{\pm 1})$ is the relative variance for $\sigma\pm 1$.
Asymmetric in potentials. Let the stiffness be $k+\alpha$ for forward beating ($\sigma=+1$) and $k-\alpha$ for reverse beating ($\sigma=-1$), with $\alpha<k$ quantifying the asymmetry ($\alpha=0$ corresponds to the symmetric case). The drag coefficient $\gamma$ is assumed to be identical for both branches. Using the SNA approach described above, the average period and its fluctuations are given by,
$\langle\mathcal{T}\rangle_{\alpha}$ $=\dfrac{1}{\left(1-\frac{\alpha^{2}}{k^{2}}\right)}\langle\mathcal{T}\rangle_{0},\text{ and }$ (22a)
$CV_{\alpha}^{2}(\mathcal{T})$ $=\dfrac{\left(1+\frac{3\alpha^{2}}{k^{2}}\right)}{\left(1-\frac{\alpha^{2}}{k^{2}}\right)}CV_{0}^{2}(\mathcal{T}),$ (22b)
where $\langle\mathcal{T}\rangle_{0}$ and $CV_{0}^{2}(\mathcal{T})$, the mean time period and relative variance for $\alpha=0$ with
$\langle\mathcal{T}\rangle_{0}{=}2\log\left(\frac{1{+}\mathcal{A}_{s}}{1{-}\mathcal{A}_{s}}\right),\text{ and }CV_{0}^{2}(\mathcal{T}){=}\frac{1}{2\epsilon g(\mathcal{A}_{s})}.$
Note that the asymmetry in stiffness enhances both the time period and the relative fluctuations about it.
Asymmetric in drag forces. Here, we assume that the drag coefficients for forward and backward beating are $\gamma+\beta$ and $\gamma-\beta$, respectively, with $\beta<\gamma$ quantifying the asymmetry ($\beta=0$ corresponds to the symmetric case). The stiffness $k$ is taken to be the same for both beatings. The expressions for the average period and the associated noise then take the form:
$\langle\mathcal{T}\rangle_{\beta}{=}\langle\mathcal{T}\rangle_{0},\text{ and }CV_{\alpha}^{2}(\mathcal{T}){=}\left(1{+}\dfrac{3\beta^{2}}{\gamma^{2}}\right)CV_{0}^{2}(\mathcal{T}).$ (23)
As can be seen from the above expressions, asymmetry in the drag coefficient does not affect the average time period; however, it increases the relative fluctuations around it.
We note that the amplitude dependence of the fluctuation noise still remains governed by $g(\mathcal{A}_{s})$, however, asymmetry modifies the overall scaling through the parameters $\alpha$ and $\beta$.
V Comparison with the rotor model
@@
where $\xi(t)$ is a Gaussian white noise with $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(0)\rangle=2D\delta(t)$. In rotor synchronization studies, the force $F$ must vary with $\theta$ [82, 83, 58, 42, 44, 11, 54], but for simplicity, here we assume $F=F_{0}$, a constant. The rotor model pumps energy continuously along the trajectory, unlike the rower model, which pumps at switching points. The rotor has a single length scale $R$, while the rower model involves two: $\mathcal{A}$ and $\mu$. The energy per cycle is $2\pi F_{0}R$ for the rotor, and $2k\mu\mathcal{A}$ for the rower.
Using SNA, the mean and relative variance for the time $\mathcal{T}$ to reach $\theta=2\pi$ from $\theta=0$ are given by
@@ VI Discussions and conclusion
The precision of oscillations is typically quantified by the quality factor $Q$, the ratio of coherence time to period. As we have shown, this is equivalent to another precision quality factor $\cal Q$}, defined as the ratio of mean current to current fluctuation. This quantity is used in stochastic thermodynamics to establish a thermodynamic uncertainty relation, setting an upper limit based on dissipation. We use the SNA to express this quantity in terms of the relative variance in first passage time (FPT), $CV^{2}(\mathcal{T})$. In the rower model, the first passage time (FPT) $\mathcal{T}$ is the time for the beating to move from $\mathcal{A}$ to $\mathcal{A}$ in a harmonic potential with stiffness $k$ centered at $\mu/2$ within a viscous medium with drag coefficient $\gamma$. The SNA provides a simple formula for $CV^{2}(\mathcal{T})$, matching numerical results closely. In the absence of a closed-form expression, the exact integral form of $CV^{2}(\mathcal{T})$ provides limited analytical clarity [2, 34], whereas the approximate SNA offers more profound insights. We found that $CV^{2}(\mathcal{T})$ is a function of scaled amplitude $\mathcal{A}_{s}=2\mathcal{A}/\mu$, with a scaling factor $k\mu^{2}/D\gamma$, representing the ratio of active energy scale to effective temperature. The SNA allows us to derive a closed-form expression of the scaling function $g({\cal A}_{s})$ which reaches a maximum at ${\cal A}_{s}=0.518$, irrespective of other parameters. Our analytical expression for the ratio of precision quality factor $\cal Q$} and energy dissipation per cycle $q_{avg}$ gives a thermodynamic uncertainty relation ${\cal Q}\,(2D\gamma/q_{avg})\leq 1$.
Finally, we examined the effect of asymmetric beating on oscillation precision. We found that asymmetry lowers the overall precision relative to the symmetric case, while leaving the dependence on oscillation amplitude unchanged. The primary effect of asymmetry is to modify the overall scaling of the precision, which now depends explicitly on the asymmetry parameter.
@@
DC and SD designed and supervised the research. SG performed the analytical and numerical calculations. All authors participated in interpreting the results and contributed to writing the manuscript.
@@
The data and codes are openly available at https://github.com/guptasubhajit1995/Rower_precision.git.